# Optimal Spectral Transitions in High-Dimensional Multi-Index Models

### v1

@@ D.1 Single-index models
D.2 $\mathrm{Cov}[\mathbf{z}|y]$ jointly diagonalizable $\forall y$
In this section, we consider the ensemble of link functions $g$ corresponding to a matrix $\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]$ that can be diagonalized with respect to the same basis of eigenvectors for all $y$: there exists an orthogonal matrix ${\bm{U}}\in\mathbb{R}^{p\times p}=(\bm{u}_{1},\ldots,\bm{u}_{p})$ independent of ${\mathrm{y}}$ and real numbers $\lambda_{k}(y)$ such that
$(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})\bm{u}_{k}=\lambda_{k}(y)\bm{u}_{k},\quad k\in\llbracket p\rrbracket.$ (112)
Consider a separate link function $\overline{g}$ defined as $\overline{g}({\bm{Z}})\coloneqq g({\bm{U}}{\bm{z}}),\;\forall{\bm{z}}\in\mathbb{R}^{p}$. It is easy to show that $g$ and $\overline{g}$ share the same weak recovery properties, including the sample complexity threshold. Indeed, introducing temporarily the notations $\overline{\mathsf{Z}}(y)$, $\overline{\mathsf{P}}(y|\;\cdot\;)$, $\overline{\alpha}_{c}$, for the quantities related to $\overline{g}$, we have that
$\overline{\mathsf{Z}}(y)$ $=(2\pi)^{-p/2}\int_{\mathbb{R}^{p}}e^{-\nicefrac{{1}}{{2}}{\bm{z}}^{T}{\bm{z}}}\delta(y-g({\bm{U}}{\bm{z}})){\rm d}{\bm{z}}$ (113)
=$\stackrel{{\scriptstyle({\bm{z}}^{\prime}={\bm{U}}{\bm{z}})}}{{=}}(2\pi)^{-p/2}\int_{\mathbb{R}^{p}}e^{-\nicefrac{{1}}{{2}}{\bm{z}}^{\prime T}{\bm{U}}{\bm{U}}^{T}{\bm{z}}^{\prime}}\delta(y-g({\bm{U}}{\bm{z}}^{\prime})){\rm d}{\bm{z}}^{\prime}=\mathsf{Z}(y),$ (114)
and similarly
$\mathbb{E}_{\bm{\mathbf{z}}}[\bm{\mathbf{z}}\overline{\mathsf{P}}(y|\bm{\mathbf{z}})]$ $={\bm{U}}^{T}\mathbb{E}_{\bm{\mathbf{z}}}[\bm{\mathbf{z}}{\mathsf{P}}(y|\bm{\mathbf{z}})],$ (115)
[…]
If $\mathbb{E}_{\bm{\mathbf{z}}}[\bm{\mathbf{z}}{\mathsf{P}}({\mathrm{y}}|\bm{\mathbf{z}})]={\bm{0}}$ almost surely for ${\mathrm{y}}\sim\mathsf{Z}$, the same applies to $\mathbb{E}_{\bm{\mathbf{z}}}[\bm{\mathbf{z}}\overline{\mathsf{P}}({\mathrm{y}}|\bm{\mathbf{z}})]$, and both problems are non-trivial. In this case
$\overline{\alpha}_{c}^{-1}$ $=\max_{\|{\bm{M}}\|_{\mathrm{F}}=1}\|\mathbb{E}_{y\sim\overline{\mathsf{Z}}}[\left(\mathbb{E}_{\bm{\mathbf{z}}}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}\overline{\mathsf{P}}(y|\bm{\mathbf{z}})]-{\bm{I}}\right){\bm{M}}\left(\mathbb{E}_{\bm{\mathbf{z}}}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}\overline{\mathsf{P}}(y|\bm{\mathbf{z}})]-{\bm{I}}\right)]\|_{\mathrm{F}}$ (117)
$=\max_{\|{\bm{M}}\|_{\mathrm{F}}=1}\|\mathbb{E}_{y\sim\mathsf{Z}}[\left(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}}\right){\bm{U}}{\bm{M}}{\bm{U}}^{T}\left(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}}\right)]\|_{\mathrm{F}}$ (118)
$=\max_{\|{\bm{M}}^{\prime}\|_{\mathrm{F}}=1}\|\mathbb{E}_{y\sim\mathsf{Z}}[\left(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}}\right){\bm{M}}^{\prime}\left(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}}\right)]\|_{\mathrm{F}}=\alpha_{c}^{-1}.$ (119)
Therefore, whenever we are in this setting, we can consider without loss of generality $(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})=\operatorname{diag}(\lambda_{1}(y),\ldots,\lambda_{p}(y))$. The eigenpairs of the operator ${\mathcal{F}}$ eq. (77) are given by
@@
with the critical sample complexity $\alpha_{c}^{-1}=\nu_{1}\coloneqq\max_{k\in\llbracket p\rrbracket}\nu_{(k,k)}$  [^3] Note that $\forall k,h\in\llbracket p^{2}\rrbracket$, $\nu_{(k,h)}\leq\max(\nu_{(k,k)},\nu_{(h,h)})$.. If the maximum is achieved by more than one pair of indices, we expect that the matrix ${\underline{\bm{L}}}$ principal eigenvalue is degenerate, with degeneracy given by the cardinality of the set =${\mathcal{I}}=\{(k,h)\in\{1,...,p\}^{2}|\nu_{(k,h)}=\alpha_{c}^{-1}\}$. Note that $\forall k,h$, $\nu_{(k,h)}=\nu_{(h,k)}$ and if $\nu_{(k,h)}=\max_{\mu\in\{k,h\}}\nu_{\mu\mu}\implies\nu_{(k,k)}=\nu_{(h,h)}$. [^4] Without loss of generality $\nu_{(h,h)}\leq\nu_{(k,k)}$ and $\nu_{(k,k)}=\nu_{(k,h)}\leq\sqrt{\nu_{(k,k)}\nu_{(h,h)}}\implies\nu_{(k,k)}\leq\nu_{(h,h)}$. Therefore $\nu_{(k,k)}=\nu_{(h,h)}$.
@@
We consider now specific cases of link functions that are such that $\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]$ is jointly diagonalizable $\forall y$. We refer to Troiani et al. [2024] for the derivation of the expressions of $\mathsf{Z}(y)$ and $\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]$ in all the examples contained in this Appendix D.2.
D.2.1 $g(z_{1},\ldots,z_{p})=p^{-1}\sum_{k\in\llbracket p\rrbracket}z_{k}^{2}$
$\mathsf{Z}(y)=\frac{e^{-\frac{py}{2}}}{y2^{p/2}\Gamma\left(\frac{p}{2}\right)}\,(p{\mathrm{y}})^{p/2},\quad\quad\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]=y{\bm{I}}.$ (125)
@@ D.2.1 $g(z_{1},\ldots,z_{p})=p^{-1}\sum_{k\in\llbracket p\rrbracket}z_{k}^{2}$
D.2.2 $g(z_{1},z_{2})=\operatorname{sign}(z_{1}z_{2})$
$\mathsf{Z}(y)=\frac{1}{2},\quad\quad\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]=\frac{2y}{\pi}\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)+{\bm{I}}.$ (129)
The matrix $(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})$ is jointly diagonalizable $\forall y$, with eigenvalues $\lambda_{1}(y)=2{\mathrm{y}}\pi^{-1}$ and $\lambda_{2}({\mathrm{y}})=-2{\mathrm{y}}\pi^{-1}$. Therefore, the eigenvalues of ${\mathcal{F}}$ are given by
@@ D.2.2 $g(z_{1},z_{2})=\operatorname{sign}(z_{1}z_{2})$
Figure 5: Distribution of the eigenvalues (dots) $\lambda\in\mathbb{C}$ of ${\underline{\bm{L}}}$ at finite $n=5\cdot 10^{3}$, for $g(z_{1},z_{2})=\operatorname{sign}(z_{1}z_{2})$, $\alpha_{c}=\nicefrac{{\pi^{2}}}{{4}}$. (Left) $\alpha=1.4<\alpha_{c}$. (Right) $\alpha=7>\alpha_{c}$. The dashed blue circle has radius equal to $\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$, i.e. the value $\gamma_{b}$ predicted in Lemma 3.4. The dashed orange vertical line corresponds to $\operatorname{Re}\lambda=\nicefrac{{\alpha}}{{\alpha_{c}}}$, the eigenvalue $\gamma_{s}$ defined in Lemma 3.3. As predicted by the state evolution equations for this problem, two significant eigenvalues (highlighted in orange) are observed near this vertical line. Additionally, one can observe that our framework predicts other two degenerate eigenvalues at $-\gamma_{s}$, here highlighted in cyan.
Figure 6: Distribution of the eigenvalues of ${\underline{\bm{T}}}$, $d=10^{4}$, for the link function $g(z_{1},z_{2})=\operatorname{sign}(z_{1}z_{2})$. The critical threshold in $\alpha_{c}=\nicefrac{{\pi^{2}}}{{4}}$. The distribution is truncated on the left. (Left) $\alpha=\alpha_{c}$. (Right) $\alpha=7>\alpha_{c}$. As predicted by the state evolution framework, in this regime we observe two eigenvalues separated from the main bulk, centered around $\lambda_{s}$ (green vertical line) obtained in Lemma 3.8. The vertical purple line correspond to the value $\lambda_{b}$ provided in Lemma 3.9 as a bound for the bulk.
D.2.3 $g(z_{1},\ldots,z_{p})=\prod_{k=1}^{p}z_{k}$
@@
$\mathsf{Z}(y)=\frac{K_{0}(|y|)}{\pi},\quad\quad\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]=\left(\begin{array}[]{cc}|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}&{\mathrm{y}}\\ {\mathrm{y}}&|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}\end{array}\right),$ (133)
@@
$\mathsf{Z}(y)=\frac{1}{(2\pi)^{p/2}}G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0\end{array}\right),\quad\quad\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]=(\lambda(y)+1){\bm{I}},$ (137)
@@
D.3 A non-jointly diagonalizable case: $g(z_{1},z_{2})=z_{1}/z_{2}$
If the matrix ${\bm{G}}(y)$ is not jointly diagonalizable $\forall y$, there is not a general simplification for equations (75,76), and each example needs to be treated separately.
@@ D.3 A non-jointly diagonalizable case: $g(z_{1},z_{2})=z_{1}/z_{2}$
$\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]$ $=\frac{1}{2\pi\mathsf{Z}(y)}\int_{\mathbb{R}^{2}}{\bm{z}}{\bm{z}}^{T}e^{-\nicefrac{{1}}{{2}}(z_{1}^{2}+z_{2}^{2})}\delta\left({\mathrm{y}}-\frac{z_{1}}{z_{2}}\right){\rm d}z_{1}{\rm d}z_{2}$ (147)
@@
The eigenpairs of $(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})$ are $\lambda_{1}(y)=1$, with eigenvector $({y},1)^{T}$, and $\lambda_{2}({y})=-1$ with eigenvector $(-1,{y})^{T}$, which depends on $y$. Considering a generic $\bm{M}=\left(\begin{array}[]{cc}m_{1}&m_{2}\\ m_{3}&m_{4}\end{array}\right)$, we have that
@@
Appendix E Details on examples - Symmetric spectral method
In all the considered examples with $p\geq 2$, the matrix $\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]$ admits a unique orthonormal basis of eigenvectors independent of $y$. Therefore, the state evolution equations can be significantly simplified following the same considerations applied in Appendix D.2, to which we refer for the notation adopted in this appendix. Additionally, for all these examples, the eigenvalues $\lambda_{k}(y)$ of $(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})$ satisfy the additional conditions
@@ Appendix E Details on examples - Symmetric spectral method
Additionally
$\forall k,h\in\llbracket p\rrbracket,\;\;\mathbb{E}_{{\mathrm{y}}}[\lambda_{k}({\mathrm{y}})\lambda_{h}({\mathrm{y}})]\geq\mathbb{E}_{{\mathrm{y}}}[\lambda_{k}({\mathrm{y}})^{2}]=\mathbb{E}_{{\mathrm{y}}}[\lambda_{h}({\mathrm{y}})^{2}].$ (165)
It is easy to verify that these conditions implies that ${\bm{V}}={\bm{I}}$, and the state evolution admits stable fixed point ${\bm{M}},{\bm{Q}}\propto{\bm{I}}$, where the proportionality constants can be numerically computed through one-dimensional integrals, expressed in terms of $\lambda_{1}(y)$ (the choice of the eigenvalue is arbitrary in this setting) and given in eq. (54).
Appendix F Proof of Proposition 3.11
@@
    By definition ${\underline{\bm{L}}}\cdot\bm{\Omega}=\gamma_{{\underline{\bm{L}}}}\bm{\Omega}$. Applying on the right of both sides $\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot((\gamma_{{\underline{\bm{L}}}}{\underline{\bm{I}}}+{\underline{\bm{G}}})^{-1}\;\cdot\;)$
    $\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot((\gamma_{{\underline{\bm{L}}}}{\underline{\bm{I}}}+{\underline{\bm{G}}})^{-1}\cdot({\underline{\bm{L}}}\cdot\bm{\Omega}))=\gamma_{{\underline{\bm{L}}}}\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot((\gamma_{{\underline{\bm{L}}}}{\underline{\bm{I}}}+{\underline{\bm{G}}})^{-1}\cdot\bm{\Omega}).$ (166)
    Recalling the definition of ${\underline{\bm{L}}}$ (16)
    $\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot(({\underline{\bm{I}}}+{\underline{\bm{G}}})^{-1}\cdot(\bm{\mathbf{X}}\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\bm{\Omega})=\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\bm{\Omega}$ (167)
    $\implies$ $\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot(({\underline{\bm{G}}}+\gamma_{{\underline{\bm{L}}}}{\underline{\bm{I}}})^{-1}\cdot(\bm{\mathbf{X}}{\bm{W}}))={\bm{W}}$ (168)
@@
    Defining $\bm{\Omega}:=({\underline{\bm{I}}}+{\underline{\bm{G}}})^{-1}\cdot(\bm{\mathbf{X}}{\bm{W}})$, we have that
    ${\underline{\bm{L}}}\cdot\bm{\Omega}$ $=(\bm{\mathbf{X}}\bm{\mathbf{X}}^{T}-\mathbb{I}_{n}){\underline{\bm{G}}}\cdot(({\underline{\bm{I}}}+{\underline{\bm{G}}}^{-1}\cdot(\bm{\mathbf{X}}{\bm{W}}))$ (169)
    $=\bm{\mathbf{X}}{\underline{\bm{T}}}\cdot{\bm{W}}-{\underline{\bm{G}}}\bm{\Omega}$ (170)
    $=\gamma_{{\underline{\bm{T}}}}({\underline{\bm{I}}}+{\underline{\bm{G}}})\cdot\bm{\Omega}-{\underline{\bm{G}}}\cdot\bm{\Omega}$ (171)
    $=\gamma_{{\underline{\bm{T}}}}\bm{\Omega}+(\gamma_{{\underline{\bm{T}}}}-1){\underline{\bm{G}}}\cdot\bm{\Omega}.$ (172)

### v2

@@ E.1 Single-index models
E.2 G$(y)$ jointly diagonalizable $\forall y$
Following the arguments at the beginning of Appendix C.3, whenever we are in this setting, we can consider without loss of generality ${\bm{G}}(y)=\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}=y]=\operatorname{diag}(\lambda_{1}(y),\ldots,\lambda_{p}(y))$. The eigenpairs of the operator ${\mathcal{F}}$ eq. (60) are given by
@@ E.2 G$(y)$ jointly diagonalizable $\forall y$
with the critical sample complexity $\alpha_{c}^{-1}=\nu_{1}\coloneqq\max_{k\in\llbracket p\rrbracket}\nu_{(k,k)}$  [^3] Note that $\forall k,h\in\llbracket p^{2}\rrbracket$, $\nu_{(k,h)}\leq\max(\nu_{(k,k)},\nu_{(h,h)})$.. If the maximum is achieved by more than one pair of indices, we expect that the matrix $\bm{L}$ principal eigenvalue is degenerate, with degeneracy given by the cardinality of the set =${\mathcal{I}}=\{(k,h)\in\{1,...,p\}^{2}|\nu_{(k,h)}=\alpha_{c}^{-1}\}$. Note that $\forall k,h$, $\nu_{(k,h)}=\nu_{(h,k)}$ and if $\nu_{(k,h)}=\max_{\mu\in\{k,h\}}\nu_{\mu\mu}\implies\nu_{(k,k)}=\nu_{(h,h)}$. [^4] Without loss of generality $\nu_{(h,h)}\leq\nu_{(k,k)}$ and $\nu_{(k,k)}=\nu_{(k,h)}\leq\sqrt{\nu_{(k,k)}\nu_{(h,h)}}\implies\nu_{(k,k)}\leq\nu_{(h,h)}$. Therefore $\nu_{(k,k)}=\nu_{(h,h)}$.
@@
We consider now specific cases of link functions that are such that $\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]$ is jointly diagonalizable $\forall y$. We refer to Troiani et al. [2025] for the derivation of the expressions of $\mathsf{Z}(y)$ and $\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]$ in all the examples contained in this Appendix E.2.
E.2.1 $g(z_{1},\ldots,z_{p})=p^{-1}\sum_{k\in\llbracket p\rrbracket}z_{k}^{2}$
$\mathsf{Z}(y)=\frac{e^{-\frac{py}{2}}}{y2^{p/2}\Gamma\left(\frac{p}{2}\right)}\,(p{\mathrm{y}})^{p/2},\quad\quad\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}|{\mathrm{y}}]=y{\bm{I}}.$ (147)
@@ E.2.1 $g(z_{1},\ldots,z_{p})=p^{-1}\sum_{k\in\llbracket p\rrbracket}z_{k}^{2}$
E.2.2 $g(z_{1},z_{2})=\operatorname{sign}(z_{1}z_{2})$
$\mathsf{Z}(y)=\frac{1}{2},\quad\quad\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}|{\mathrm{y}}]=\frac{2y}{\pi}\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)+{\bm{I}}.$ (151)
The matrix $\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]$ is jointly diagonalizable $\forall y$, with eigenvalues $\lambda_{1}(y)=2{\mathrm{y}}\pi^{-1}$ and $\lambda_{2}({\mathrm{y}})=-2{\mathrm{y}}\pi^{-1}$. Therefore, the eigenvalues of ${\mathcal{F}}$ are given by
@@ E.2.2 $g(z_{1},z_{2})=\operatorname{sign}(z_{1}z_{2})$
Figure 5: Distribution of the eigenvalues (dots) $\lambda\in\mathbb{C}$ of $\bm{L}$ at finite $n=5\cdot 10^{3}$, for $g(z_{1},z_{2})=\operatorname{sign}(z_{1}z_{2})$, $\alpha_{c}=\nicefrac{{\pi^{2}}}{{4}}$. (Left) $\alpha=1.4<\alpha_{c}$. (Right) $\alpha=7>\alpha_{c}$. The dashed blue circle has radius equal to $\sqrt{\nicefrac{{\alpha}}{{\alpha_{c}}}}$, i.e. the value $\gamma_{b}$ predicted in Theorem 3.4. The dashed orange vertical line corresponds to $\operatorname{Re}\lambda=\nicefrac{{\alpha}}{{\alpha_{c}}}$, the eigenvalue $\gamma_{s}$ defined in Theorem 3.3. As predicted by the state evolution equations for this problem, two significant eigenvalues (highlighted in orange) are observed near this vertical line. Additionally, one can observe that our framework predicts other two degenerate eigenvalues at $-\gamma_{s}$, here highlighted in cyan.
Figure 6: Distribution of the eigenvalues of $\bm{T}$, $d=10^{4}$, for the link function $g(z_{1},z_{2})=\operatorname{sign}(z_{1}z_{2})$. The critical threshold in $\alpha_{c}=\nicefrac{{\pi^{2}}}{{4}}$. The distribution is truncated on the left. (Left) $\alpha=\alpha_{c}$. (Right) $\alpha=7>\alpha_{c}$. As predicted by the state evolution framework, in this regime we observe two eigenvalues separated from the main bulk, centered around $\lambda_{s}$ (green vertical line) obtained in Theorem 3.8. The vertical purple line correspond to the value $\lambda_{b}$ provided in Theorem 3.9 as a bound for the bulk.
E.2.3 $g(z_{1},\ldots,z_{p})=\prod_{k=1}^{p}z_{k}$
@@
$\mathsf{Z}(y)=\frac{K_{0}(|y|)}{\pi},\quad\quad\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}|{\mathrm{y}}]=\left(\begin{array}[]{cc}|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}&{\mathrm{y}}\\ {\mathrm{y}}&|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}\end{array}\right),$ (155)
@@
$\mathsf{Z}(y)=\frac{1}{(2\pi)^{p/2}}G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0\end{array}\right),\quad\quad\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}|{\mathrm{y}}]=(\lambda(y)+1){\bm{I}},$ (159)
@@
E.3 A non-jointly diagonalizable case: $g(z_{1},z_{2})=z_{1}/z_{2}$
If the matrix ${\bm{G}}(y)$ is not jointly diagonalizable $\forall y$, there is not a general simplification for equations (58,59), and each example needs to be treated separately.
@@ E.3 A non-jointly diagonalizable case: $g(z_{1},z_{2})=z_{1}/z_{2}$
$\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}|{\mathrm{y}}]$ $=\frac{1}{2\pi\mathsf{Z}(y)}\int_{\mathbb{R}^{2}}{\bm{z}}{\bm{z}}^{T}e^{-\nicefrac{{1}}{{2}}(z_{1}^{2}+z_{2}^{2})}\delta\left({\mathrm{y}}-\frac{z_{1}}{z_{2}}\right){\rm d}z_{1}{\rm d}z_{2}$ (169)
@@
The eigenpairs of ${\bm{G}}(y)$ are $\lambda_{1}(y)=1$, with eigenvector $({y},1)^{T}$, and $\lambda_{2}({y})=-1$ with eigenvector $(-1,{y})^{T}$, which depends on $y$. Considering a generic $\bm{M}=\left(\begin{array}[]{cc}m_{1}&m_{2}\\ m_{3}&m_{4}\end{array}\right)$, we have that
@@
Appendix F Details on examples - Symmetric spectral method
In all the considered examples with $p\geq 2$, the matrix $\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}|{\mathrm{y}}=y]$ admits a unique orthonormal basis of eigenvectors independent of $y$. Therefore, the state evolution equations can be significantly simplified following the same considerations applied in Appendices C.3 and E.2, to which we refer for the notation adopted in this appendix. Additionally, for all these examples, the eigenvalues $\lambda_{k}(y)$ of ${\bm{G}}(y)$ satisfy the additional conditions
@@ Appendix F Details on examples - Symmetric spectral method
so that
$\forall k,h\in\llbracket p\rrbracket,\;\;\mathbb{E}_{{\mathrm{y}}}[\lambda_{k}({\mathrm{y}})^{2}]=\mathbb{E}_{{\mathrm{y}}}[\lambda_{h}({\mathrm{y}})^{2}]\geq\mathbb{E}_{{\mathrm{y}}}[\lambda_{k}({\mathrm{y}})\lambda_{h}({\mathrm{y}})].$ (187)
It is easy to verify that these conditions implies that ${\bm{V}}={\bm{I}}$, and the state evolution admits stable fixed point ${\bm{M}},{\bm{Q}}\propto{\bm{I}}$, where the proportionality constants can be numerically computed through one-dimensional integrals, expressed in terms of $\lambda_{1}(y)$ (the choice of the eigenvalue is arbitrary in this setting) and given in eq. (36). Additionally, the largest eigenvalue of ${\bm{T}}$ is always equal to one and degenerate (see App. C.3).
Proposition 3.12 readily implies that, for all these examples, the matrix ${\bm{L}}$ has a correspondent informative subspace of eigenvectors that can be computed from the subspace of leading eigenvectors of ${\bm{T}}$ with eigenvalue equal to 1 and hidden in the bulk.
Appendix G Proof of Proposition 3.12
@@
    By definition $\bm{L}\bm{\omega}=\gamma_{\bm{L}}\bm{\omega}$. Applying on the right of both sides $\hat{\bm{G}}(\gamma_{\bm{L}}{\bm{I}}_{np}+\hat{\bm{G}})^{-1}$
    $\hat{\bm{G}}(\gamma_{\bm{L}}{\bm{I}}_{np}+\hat{\bm{G}})^{-1}\bm{L}\bm{\omega}=\gamma_{\bm{L}}\hat{\bm{G}}(\gamma_{\bm{L}}{\bm{I}}_{np}+\hat{\bm{G}})^{-1}\bm{\omega}.$ (188)
    Recalling the definition of $\bm{L}$ (11), for $i\in\llbracket n\rrbracket,\,\mu\in\llbracket p\rrbracket,\,k\in\llbracket d\rrbracket$
    $\sum_{j,\ell\in\llbracket n\rrbracket}\sum_{\nu\in\llbracket p\rrbracket}\left[\hat{{\bm{G}}}\left(\gamma_{{\bm{L}}}{\bm{I}}_{np}+\hat{{\bm{G}}}\right)^{-1}\right]_{(i\mu),(j\nu)}\left[\bm{\mathbf{X}}\bm{\mathbf{X}}^{T}\right]_{j\ell}\,\left[\hat{{\bm{G}}}\bm{\omega}\right]_{(\ell\nu)}=\left[\hat{{\bm{G}}}\bm{\omega}\right]_{(i\mu)}$ (189)
    $\implies$ $\sum_{\nu\in\llbracket p\rrbracket}\left[{\bm{G}}({\mathrm{y}}_{i})\left(\gamma_{{\bm{L}}}{\bm{I}}_{p}+{\bm{G}}({\mathrm{y}}_{i})\right)^{-1}\right]\,\sum_{h\in\llbracket d\rrbracket}{\mathrm{X}}_{ih}\underbrace{\sum_{j\in\llbracket n\rrbracket}{\mathrm{X}}_{jh}\left[\hat{{\bm{G}}}\bm{\omega}\right]_{(j\nu)}}_{=w_{(h\nu)}}=\left[\hat{{\bm{G}}}\bm{\omega}\right]_{(i\mu)}$ (190)
    $\stackrel{{\scriptstyle\bm{\mathbf{X}}^{T}\,\cdot}}{{\implies}}$ $\sum_{i\in\llbracket n\rrbracket}X_{ik}\sum_{\nu\in\llbracket p\rrbracket}\left[{\bm{G}}({\mathrm{y}}_{i})\left(\gamma_{{\bm{L}}}{\bm{I}}_{p}+{\bm{G}}({\mathrm{y}}_{i})\right)^{-1}\right]\,\sum_{h\in\llbracket d\rrbracket}{\mathrm{X}}_{ih}w_{(h\nu)}=w_{(i\mu)}$ (191)
    $\implies$ ${\bm{T}}_{\gamma_{{\bm{L}}}}{\bm{w}}={\bm{w}}$ (192)
@@
    Defining $\bm{\omega}:=({\bm{I}}_{np}+\hat{\bm{G}})^{-1}{\rm vec}\left({\bm{X}}{\rm mat}({\bm{w}})\right)$, we have that, for $i\in\llbracket n\rrbracket,\mu\in\llbracket n\rrbracket$
    $[{\bm{L}}\bm{\omega}]_{(i\mu)}$ $=\sum_{j\in\llbracket n\rrbracket}\sum_{\nu\in\llbracket p\rrbracket}([\bm{\mathbf{X}}\bm{\mathbf{X}}^{T}]_{ij}-\delta_{ij})\left[\bm{G}({\mathrm{y}}_{j})\left({\bm{I}}+\bm{G}({\mathrm{y}}_{j})\right)^{-1}]\right]_{\mu\nu}\sum_{h\in\llbracket d\rrbracket}{\mathrm{X}}_{jh}w_{(h\nu)}$ (193)
    $=\sum_{h\in\llbracket d\rrbracket}{\mathrm{X}}_{ih}[{\bm{T}}{\bm{w}}]_{(h\mu)}-[\hat{{\bm{G}}}\bm{\omega}]_{i\mu}$ (194)
    $=\left[\left(\gamma_{{\bm{T}}}({\bm{I}}_{np}+\hat{{\bm{G}}})-\hat{{\bm{G}}}\right)\bm{\omega}\right]_{(i\mu)}$ (195)